Operations Report – February 12th

Crew 292 Operations Report 12-02-2024

SOL: #8

Name of person filing report: Rajvi Patel

Non-nominal systems: None

Notes on non-nominal systems: None


Spirit rover used: EVA #9

Hours: (before EVA): 245.8

Beginning charge: (Before EVA): 86 %

Ending charge: (On return from EVA, before recharging): 47%

Currently charging: No

Opportunity rover used: EVA #10

Hours: 161.5

Beginning charge: 100

Ending charge: 63

Currently charging: No

Curiosity rover used: EVA #10

Hours: 265.5

Beginning charge: 100

Ending charge: 58%

Currently charging: No

Perseverance rover used: EVA #9

Hours: 283.9

Beginning charge: 81%

Ending charge: 48%

Currently charging: No

General notes on rovers:

Summary of Hab operations: put notes here

WATER USE: 25.73 gallons

Water (static tank): 284.99 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of Internet: Nominal

Summary of suits and radios: Suits and radios worked fine for today’s EVAs

Status on suit #6: I have replaced the battery and have completed the 1-hour test. It was 13.0 V and reduced to 11.9 V after 1 hour. Suit #6 is now operational.

Summary of GreenHab operations: put notes here

WATER USE: 7.2 gallons

Heater: On

Supplemental light: 22:00-02:00

Harvest: 4g basil, 30 g lettuce, 47g tomatoes, 60g kale, 36g Radish, 20 g Cilantro.

Summary of ScienceDome operations: On EVA 9 in the morning two samples were taken, one from the soil in a crevice worn out of a larger rock, the other from a yellow goo (crew journalist’s description). In the afternoon DNA extraction was performed on the collected samples and non-targeted PCR was run on the previously extracted samples from yesterday; two water samples and two soil samples. As this report is being written the gel to visualize results is running. Our Crew Journalist Clare used it to clean rock specimens.

Dual split: Heat On

Summary of RAM operations: (Tools used, work done) Made necessary arrangements for a fire extinguisher as requested by the mission support.

Summary of any observatory issues: put notes here: None

Summary of health and safety issues: put notes here: None

Questions, concerns and requests to Mission Support: put notes here: None

Copyright © The Mars Society. All rights reserved. | Main Site